OpenStudy (anonymous):

Raven’s parents bought a 24’ sailboat boat for $32,000. After 2 years, the boat was worth $28,500. Using x as the number of years and y as the boat value, graph the two points. Part 1: [2 points] Describe the slope of the line which connects the two points. Part 2: [3 points] How can your graph be used to predict the value of the boat in 5 years? Part 3: [2 points] What is the projected boat value in 5 years?

jimthompson5910 (jim_thompson5910):

Part 1 Use the slope formula m = (y2-y1)/(x2-x1) m = (28500-32000)/(2-0) m = ???

OpenStudy (anonymous):

so what do you think it is?

jimthompson5910 (jim_thompson5910):

28500-32000 = ???

OpenStudy (anonymous):

so do i put 28500-32000?

jimthompson5910 (jim_thompson5910):

no I'm asking you what is the result of that subtraction

OpenStudy (anonymous):

-3500?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

2-0 = 2

jimthompson5910 (jim_thompson5910):

so m = (28500-32000)/(2-0) turns into m = (-3500)/(2) which becomes m = -1750

jimthompson5910 (jim_thompson5910):

so the slope is -1750 this means that the value of the boat is decreasing $1750 each year

jimthompson5910 (jim_thompson5910):

for part 2, you just graph the two points (0,32000) and (2,28500) and draw a straight line through them this will allow you to estimate the value of the boat in 5 years

OpenStudy (anonymous):

$8750 for 5 years

jimthompson5910 (jim_thompson5910):

optionally, you can use the points (0,32) and (2,28.5) and it will be implied that y is the value in thousands of dollars

OpenStudy (anonymous):

$32000-$8750?